# Self-supervised pretraining of vision transformers for animal behavioral analysis and neural encoding

### v1

@@ Results
Beast demonstrates strong action segmentation performance across all datasets (Fig. 4). Remarkably, ImageNet-pretrained ViT-M patch embeddings nearly match keypoint-based methods despite utilizing a frozen, general-purpose backbone. This establishes a competitive baseline without requiring the thousands of labeled frames needed to train pose estimation networks. On the IBL dataset, Beast improves upon ImageNet baselines. The keypoint-based model excels here due to action classes corresponding to single paw movements that are easily captured by pose estimation, but this advantage disappears with ensembling: Beast ensemble performance matches the keypoint ensemble. The CalMS21 dataset better demonstrates Beast’s capabilities, where our method surpasses the SimBA baseline and substantially outperforms the TREBA baseline (Table 2). Additional experiments confirm domain-specific pretraining benefits: Beast CLS tokens consistently outperform their ImageNet-pretrained counterparts (Table 2), though patch-based models perform significantly better due to their enhanced spatial resolution and multi-headed attention pooling. Across all experiments, nonlinear models consistently outperform their linear counterparts, except for PCA features on CalMS21 (Table 13). All evaluations use frozen backbones with only linear/TCN heads fine-tuned, suggesting further gains may be possible through full backbone fine-tuning.
Table 2: Action segmentation performance (F1$\pm$±S.E.M.).
@@
Beast’s advantages extend beyond absolute F1 improvements. Pose estimation-based approaches require both extensive labeling and iterative training and validation of pose estimation models before action segmentation, often a months- or even years-long process [30]. Beast eliminates this entire pipeline, achieving competitive or superior performance using only unlabeled video for pretraining.
@@ 5 Discussion
This work introduces Beast, a framework for self-supervised vision transformer pretraining leveraging domain-specific video data. We demonstrated Beast’s significant benefits across neural encoding, pose estimation, and action segmentation tasks. While Beast requires substantial computational resources (hundreds of GPU hours for large-scale pretraining on a single dataset), this initial investment creates a versatile backbone applicable to multiple downstream tasks. For many laboratories, this computational cost is far outweighed by the dramatic reduction in human labeling time and the improved accuracy achieved without iterative model refinement. The black-box nature of ViT embeddings presents interpretability challenges in scientific contexts where transparent representations like pose estimates are often preferred. Future work will focus on model compression strategies to reduce training time without performance degradation, and further exploring the transformer attention mechanism to better understand how Beast functions in different task contexts.
@@
This work was supported by the following: Gatsby Charitable Foundation GAT3708, NIH U19NS123716, NSF 1707398, the National Science Foundation and DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence), Simons Foundation, Wellcome Trust 216324, and funds from Zuckerman Institute Team Science.
@@ A.1 IBL
This dataset [26] from the International Brain Lab (IBL) and consists of head-fixed mice performing a decision-making task [27, 28, 29]. Two cameras–‘left’ (60 Hz) and ‘right’ (150 Hz)–capture roughly orthogonal side views of the mouse’s face and upper trunk during each session. Frames are downsampled to 256 $\times$ 320 pixels for labeling and video storage. We accessed the raw videos and neural activity under the CC-BY 4.0 license using these instructions: https://int-brain-lab.github.io/iblenv/notebooks_external/data_release_brainwidemap.html. The data can also be visualized through a browser at https://viz.internationalbrainlab.org.
Pose estimation
Frames were reshaped during training to 256 $\times$ 256 pixels. Two keypoints were labeled per view, one for each paw. We accessed the initial pose estimation labels from the public repository at https://ibl-brain-wide-map-public.s3.amazonaws.com/aggregates/Tags/2023_Q1_Biderman_Whiteway_et_al/_ibl_videoTracking.trainingDataPaw.7e79e865-f2fc-4709-b203-77dbdac6461f.zip under the CC-BY 4.0 license. This dataset contains 6,071 labeled train frames from 35 animals and 1,446 labeled test frames from 10 animals, all at 128 $\times$ 102 pixel resolution.
Despite the large number of labeled frames, we observed poor performance in sessions with bright lights or other unusual distractors. Additionally, the low resolution often obscured fine details, for example making it difficult to visually distinguish individual paws when they are close together. To address these limitations, we retrieved the full-resolution frames (1280 $\times$ 1024 for left view, 640 $\times$ 512 for right) from the raw videos and downsampled them to 320 $\times$ 256 pixels. This higher resolution revealed occasional labeling errors, which we manually corrected. We then added 1,437 newly labeled frames from 15 additional animals, creating an expanded training set of 7,609 frames from 50 animals. We will publicly release these updated labels under a CC-BY 4.0 license upon acceptance of this manuscript.
Action segmentation
@@
We expanded this dataset as the original study [4] only used five animals each for training and testing. First, we trained an ensemble of five TCN-based action segmentation models on all 10 existing animals. We applied these models to a new batch of 53 sessions and calculated the variance in predicted probabilities across all models for each frame. The 19 sessions with the highest average ensemble variance (indicating where models disagreed most) were selected for further labeling. We then labeled an additional 36,009 frames from these sessions. For the analyses in this paper, we selected the subset of all labeled sessions that are included in the Beast pretraining sessions, and split these into train (32,521 frames from 18 animals) and test sets (7,786 frames from 5 animals). We will publicly release these updated labels under a CC-BY 4.0 license upon acceptance of this manuscript.
Neural encoding
For the neural analysis we use a subset of the IBL repeated site dataset [28]. This dataset consists of Neuropixels recordings collected from 10 labs with standardized experimental pipelines. The recordings target the same five brain regions across all mice: VISa (primary visual cortex), CA1 and DG (hippocampus), and LP and PO (thalamic nuclei). We evaluate neural encoding models on five randomly selected sessions. Moreover, we used the trial-aligned neural activity data, taking 2 seconds of activity aligned to wheel movement onset. We binned the spikes every 20 ms to get a total of 100 bins per trial. We also filtered out the low firing rate neurons by setting a minimum threshold of 2 Hz.
@@ A.2 IBL-whisker
We localize the whisker pad using anchor keypoints on the nose and eye, following the procedure in [30]. We use the same sessions and neural activity as the “IBL” dataset.
@@ A.3 Mirror-mouse
Head-fixed mice ran on a circular treadmill while avoiding a moving obstacle [72]. The treadmill had a transparent floor and a mirror mounted inside at 45°, allowing a single camera to capture two roughly orthogonal views (side view and bottom view via the mirror) at 250 Hz. The camera was positioned at a large distance from the subject ($\sim$1.1 m) to minimize perspective distortion. Frames are 406 $\times$ 396 pixels and reshaped during pose estimation training to 256 $\times$ 256 pixels. Seventeen keypoints were labeled across the two views including seven keypoints on the mouse’s body per view, plus three keypoints on the moving obstacle. The full training dataset consists of 789 labeled frames across 10 animals; the test dataset consists of 253 labeled frames across three animals. We accessed the labeled pose estimation dataset from https://doi.org/10.6084/m9.figshare.24993315.v1 under the CC-BY 4.0 license.
@@ A.4 Mirror-fish
Mormyrid fish of the species Gnathonemus petersii swam freely in and out of an experimental tank, capturing worms from a well [3]. The tank had a side mirror and a top mirror, both at 45°, providing three different views seen from a single camera at 300 Hz. The camera was placed $\sim$1.7 m away from the center of the fish tank to reduce distortions. Frame are 384 $\times$ 512 pixels and reshaped during training to 256 $\times$ 384 pixels. Seventeen body parts were labeled across each of three views for a total of 51 keypoints. The full training dataset consists of 373 frames across three animals; the test dataset consists of 94 frames across three animals. We accessed the labeled pose estimation dataset from https://doi.org/10.6084/m9.figshare.24993363.v1 under the CC-BY 4.0 license.
@@ A.5 CRIM13
The Caltech Resident-Intruder Mouse (CRIM13) dataset [6] consists of two mice interacting in an enclosed arena, captured by top and side-view cameras at 30 Hz. We only used the top view. Frames are 480 $\times$ 640 pixels and reshaped during training to 256 $\times$ 256 pixels. Seven keypoints were labeled on each mouse for a total of 14 keypoints [57]. The full training dataset consists of 3,986 frames across four resident mice; the test dataset consists of 1,274 frames across the same four resident mice but a different set of intruder mice. The original dataset is available at https://data.caltech.edu/records/4emt5-b0t10. We accessed the labeled pose estimation dataset from https://doi.org/10.6084/m9.figshare.24993384.v1 under the CC-BY 4.0 license.
@@ A.6 CalMS21
The Caltech Mouse Social Interactions (CalMS21) dataset [59], like CRIM13, consists of two mice interacting in an enclosed arena, captured by a top-view camera at 30 Hz. The dataset consists of many videos with tracked poses and corresponding frame-level behavior annotations. Four behavior classes are labeled: attack, investigation, mount, and other (i.e., none of the above). The full training dataset consists of 506,668 frames across 68 videos; the test dataset consists of 262,107 frames across 19 videos. We accessed the pose estimates, TREBA features [60], and behavior annotations from https://doi.org/10.22002/D1.1991 under the CC-BY 4.0 license.
@@ A.7 Facemap
Head-fixed mice were free to run on an air-floating ball in darkness [63]. A single infrared camera captured one of several side or front views of the mouse’s face and upper trunk during each session at 50 Hz. Fifteen keypoints were labeled across the face (mouse, nose, whiskers, eyes) and paw. Neural activity was recorded across visual and sensorimotor areas using two-photon calcium imaging. Approximately 30,000 to 50,000 cells were recorded in a given session. In our encoding task, we predict the 128 neural principal components following [63]. We evaluate neural encoding models on five randomly selected sessions. The publicly available data did not contain additional videos, so we only fine-tuned neural encoding models with this dataset. We accessed the raw videos, pose estimates, and neural activity from https://doi.org/10.25378/janelia.23712957 under the CC-BY-NC 4.0 license.
Table 3: Number of training/validation/test frames utilized across tasks, with number of source videos in parentheses. Pretraining frames are unlabeled. Pose estimation and action segmentation frames are labeled; these models are trained across multiple videos. Neural encoding frames have matched neural activity (output) for each time point of behavior (input); these models are trained on single videos, since the neural populations change from one session to the next.
Pretraining Pose estimation Action segmentation Neural encoding train/val test train/val test train/val test IBL [29] 138,600 (77) 7,609 (128) 1,446 (19) 35,521 (18) 7,786 (5) 338,760 (5) 42,720 (5) Mirror-mouse [72] 94,252 (17) 789 (17) 253 (5) - - - - Mirror-fish [3] 47,921 (28) 373 (28) 94 (10) - - - - CRIM13 [6] 99,914 (37) 3,986 (37) 1,274 (19) - - - - CalMS21 [59] 103,544 (37) - - 506,668 (68) 262,107 (19) - - Facemap [63] - - - - - 1,790,200 (5) 447,550 (5)
@@ Appendix B Beast implementation
Beast utilizes a standard ViT-B/16 architecture [12], and combines a masked autoencoding and temporal contrastive learning loss. This approach is also taken by ViC-MAE [17], and we introduce key adaptations and simplifications to make Beast suitable for applications in behavioral neuroscience, which we elaborate on more in the following sections.

### v2

@@ Results.
Beast demonstrates strong action segmentation performance across all datasets (Fig. 4). Remarkably, ImageNet-pretrained ViT-M patch embeddings nearly match keypoint-based methods despite utilizing a frozen, general-purpose backbone. This establishes a competitive baseline without requiring the thousands of labels needed to train pose estimation networks. On IBL, Beast improves upon ImageNet baselines. The keypoint-based model excels here due to action classes corresponding to paw movements easily captured by pose estimation, but this advantage disappears with ensembling: Beast ensemble F1 matches the keypoint ensemble. CalMS21 better demonstrates Beast’s abilities, which surpasses the SimBA baseline and substantially outperforms the TREBA baseline (Table 2). The ensembled F1 score of 0.84 places our result in the top 15 of the Multi-Agent Behavior Challenge on AIcrowd.com (top score of 0.89). Additional experiments confirm domain-specific pretraining benefits: Beast CLS tokens consistently outperform their ImageNet-pretrained counterparts (Table 2), though patch-based models perform significantly better due to their enhanced spatial resolution and multi-headed attention pooling. An ablation experiment on Beast’s loss terms show that backbones pretrained with a contrastive-only (ViT-C) or MAE-only (ViT-M) loss do not perform as well as their combination (Table 15). Across all experiments, nonlinear models consistently outperform their linear counterparts, except for PCA features on CalMS21 (Table 15). All evaluations use frozen backbones with only linear/TCN heads fine-tuned, suggesting further gains may be possible through full backbone fine-tuning.
[…]
@@
Beast’s advantages extend beyond absolute F1 improvements. Pose estimation-based approaches require both extensive labeling and iterative training and validation of pose estimation models before action segmentation, often a months- or even years-long process [IBL et al., 2022]. Beast eliminates this entire pipeline, achieving competitive or superior performance using only unlabeled video for pretraining.
@@ 5 Discussion
This work introduces Beast, a framework for self-supervised vision transformer pretraining leveraging domain-specific video data. We demonstrated Beast’s significant benefits across neural encoding, pose estimation, and action segmentation tasks. Notably, Beast outperforms DINOv2 and other computer vision foundation models across all tasks (Table 13), demonstrating the power of domain-specific pretraining. Our frame-based approach is an efficient alternative to native video models like VideoMAE [Tong et al., 2022], which require significantly more compute for training and inference (Table 8); however Beast still outperforms a frozen VideoMAE on the neural encoding task (Table 13).
Our work establishes a foundation for several promising future directions. Investigation of transformer attention and learned features could clarify how Beast operates across different tasks. The black-box nature of ViT embeddings presents interpretability challenges in scientific contexts where transparent representations like pose estimates are often preferred. Visualization methods provide initial insights (Fig. 7), but systematic analysis of what features drive performance on different tasks would strengthen our understanding of when and why Beast succeeds.
While we have demonstrated Beast’s performance across diverse experimental contexts–including head-fixed and freely moving animals, single- and multi-view setups, and solitary or social behaviors–validation across more environments and species is needed. The success of masked autoencoding and contrastive losses in general computer vision suggests Beast should adapt well to naturalistic settings (e.g., home cages, zoos, field studies). The primary challenge will be adjusting the frame sampling strategy to accommodate different visual statistics and behavioral distributions in these less controlled environments.
Finally, we see two complementary paths toward making powerful self-supervised models more accessible to individual labs. First, using smaller transformer architectures will reduce pretraining, fine-tuning, and inference costs, but may sacrifice performance. Second, Beast’s framework could enable foundation models of animal behavior trained across diverse datasets, rather than the dataset-specific pretraining we present here. Such foundation models would allow labs to finetune already-powerful pretrained models rather than pretraining themselves. Together, these approaches would lower barriers to adoption and enable wider application of self-supervised learning across the neuroscience community.
@@
This work was supported by Gatsby Charitable Foundation GAT3708, NIH U19NS123716 and R50NS145433, NSF 1707398, the National Science Foundation and DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence), Simons Foundation, Wellcome Trust 216324, and funds from Zuckerman Institute Team Science.
Reproducibility statement
We have attempted to provide sufficient detail to foster reproducibility of all analyses presented in this manuscript.
Data availability.
All datasets are sourced from public repositories, with links and accompanying licenses provided in Appendix A.
Model availability.
Pretrained Beast models for each dataset are available at https://huggingface.co/collections/paninski-lab/beast under the CC-BY 4.0 license.
Code availability.
Analysis details.
  • Pretraining (Appendix B): Beast pretraining procedure and frame selection strategy.
  • Neural encoding (Appendix C): Model architectures, training procedures, and hyperparameter tuning details.
  • Pose estimation (Appendix D): Training procedures and hyperparameter selection.
  • Action segmentation (Appendix E): Model architectures, training procedures, and hyperparameter tuning details.
@@ A.1 IBL
This dataset [IBL, 2023] from the International Brain Lab (IBL) and consists of head-fixed mice performing a decision-making task [IBL et al., 2021, 2025b, 2025a]. Two cameras–‘left’ (60 Hz) and ‘right’ (150 Hz)–capture roughly orthogonal side views of the mouse’s face and upper trunk during each session. Frames are downsampled to 256 $\times$ 320 pixels for labeling and video storage. We accessed the raw videos and neural activity under the CC-BY 4.0 license using these instructions: https://int-brain-lab.github.io/iblenv/notebooks_external/data_release_brainwidemap.html. Frames for pretraining are available at https://doi.org/10.6084/m9.figshare.31310143 under the CC-BY 4.0 license.
Pose estimation.
Frames were reshaped during training to 256 $\times$ 256 pixels. Two keypoints were labeled per view, one for each paw. We accessed the initial pose estimation labels from the public repository at https://ibl-brain-wide-map-public.s3.amazonaws.com/aggregates/Tags/2023_Q1_Biderman_Whiteway_et_al/_ibl_videoTracking.trainingDataPaw.7e79e865-f2fc-4709-b203-77dbdac6461f.zip under the CC-BY 4.0 license. This dataset contains 6,071 labeled train frames from 35 animals and 1,446 labeled test frames from 10 animals, all at 128 $\times$ 102 pixel resolution. Despite the large number of labeled frames, we observed poor performance in sessions with bright lights or other distractors. Additionally, the low resolution often obscured fine details, for example complicating discrimination of individual paws when close together. To address these limitations, we retrieved the full-resolution frames (1280 $\times$ 1024 for left view, 640 $\times$ 512 for right) from the raw videos and downsampled them to 320 $\times$ 256 pixels. This higher resolution revealed occasional labeling errors, which we manually corrected. We then added 1,437 newly labeled frames from 15 additional animals, creating an expanded training set of 7,609 frames from 50 animals. The updated pose estimation dataset is available at https://ibl-brain-wide-map-public.s3.amazonaws.com/aggregates/Tags/2026_Q1_Wang_Yu_et_al/_ibl_videoTracking.trainingDataPaw.zip under the CC-BY 4.0 license.
Action segmentation.
@@
We expanded this dataset as the original study [Blau et al., 2024] only used five animals each for training and testing. First, we trained an ensemble of five TCN-based action segmentation models on all 10 existing animals. We applied these models to a new batch of 53 sessions and calculated the variance in predicted probabilities across all models for each frame. The 19 sessions with the highest average ensemble variance (indicating where models disagreed most) were selected for further labeling. We then labeled 36,009 additional frames from these sessions. For the analyses in this paper, we selected the subset of all labeled sessions that are included in the Beast pretraining sessions, and split these into train (32,521 frames from 18 animals) and test sets (7,786 frames from 5 animals). The updated action segmentation dataset is available at https://ibl-brain-wide-map-public.s3.amazonaws.com/aggregates/Tags/2026_Q1_Wang_Yu_et_al/_ibl_videoActionSegmentation.trainingDataNearPaw.zip under CC-BY 4.0 license.
Neural encoding.
For the neural analysis we use a subset of the IBL repeated site dataset [IBL et al., 2025b]. This dataset consists of Neuropixels recordings collected from 10 labs with standardized experimental pipelines. The recordings target the same five brain regions across all mice: VISa (primary visual cortex), CA1 and DG (hippocampus), and LP and PO (thalamic nuclei). We evaluate neural encoding models on five randomly selected sessions. Moreover, we used the trial-aligned neural activity data, taking 2 seconds of activity aligned to wheel movement onset. We binned the spikes every 20 ms to get a total of 100 bins per trial. We also filtered out the low firing rate neurons by setting a minimum threshold of 2 Hz. For each trial we randomly select 100 video frames (out of a possible 120) to fit the downstream neural encoding models, resulting in an effective sampling rate of 50 Hz to match the neural data.
@@ A.2 IBL-whisker
We localize the whisker pad using anchor keypoints on the nose and eye, following the procedure in IBL et al. [2022]. We use the same sessions and neural activity as the “IBL” dataset.
@@ A.3 Mirror-mouse
Head-fixed mice ran on a circular treadmill while avoiding a moving obstacle [Warren et al., 2021]. The treadmill had a transparent floor and a mirror mounted inside at 45°, allowing a single camera to capture two roughly orthogonal views (side view and bottom view via the mirror) at 250 Hz. The camera was positioned at a large distance from the subject ($\sim$1.1 m) to minimize perspective distortion. Frames are 406 $\times$ 396 pixels and reshaped during pose estimation training to 256 $\times$ 256 pixels. Seventeen keypoints were labeled across the two views including seven keypoints on the mouse’s body per view, plus three keypoints on the moving obstacle. The full training dataset consists of 789 labeled frames across 10 animals; the test dataset consists of 253 labeled frames across three animals. We accessed the labeled pose estimation dataset from https://doi.org/10.6084/m9.figshare.24993315.v1 under the CC-BY 4.0 license. Frames for pretraining are available at https://doi.org/10.6084/m9.figshare.31320529 under the CC-BY 4.0 license.
@@ A.4 Mirror-fish
Mormyrid fish of the species Gnathonemus petersii swam freely in and out of an experimental tank, capturing worms from a well [Biderman et al., 2024, Pedraja et al., 2025]. The tank had a side mirror and a top mirror, both at 45°, providing three different views seen from a single camera at 300 Hz. The camera was placed $\sim$1.7 m away from the center of the fish tank to reduce distortions. Frames are 384 $\times$ 512 pixels and reshaped during training to 256 $\times$ 384 pixels. Seventeen body parts were labeled across each of three views for a total of 51 keypoints. The full training dataset consists of 373 frames across three animals; the test dataset consists of 94 frames across three animals. We accessed the labeled pose estimation dataset from https://doi.org/10.6084/m9.figshare.24993363.v1 under the CC-BY 4.0 license. Frames for pretraining are available at https://doi.org/10.6084/m9.figshare.31320775 under the CC-BY 4.0 license.
@@ A.5 CRIM13
The Caltech Resident-Intruder Mouse (CRIM13) dataset [Burgos-Artizzu et al., 2012] consists of two mice interacting in an enclosed arena, captured by a top-view camera at 30 Hz. Frames are 480 $\times$ 640 pixels and reshaped during pose estimation training to 256 $\times$ 256 pixels. Seven keypoints were labeled on each mouse for a total of 14 keypoints [Segalin et al., 2021]. The full training dataset consists of 3,986 frames across four resident mice; the test dataset consists of 1,274 frames across the same four resident mice but a different set of intruder mice. The original dataset is available at https://data.caltech.edu/records/4emt5-b0t10. We accessed the labeled pose estimation dataset from https://doi.org/10.6084/m9.figshare.24993384.v1 under the CC-BY 4.0 license. Frames for pretraining are available at https://doi.org/10.6084/m9.figshare.31321213 under the CC-BY 4.0 license.
@@ A.6 CalMS21
The Caltech Mouse Social Interactions (CalMS21) dataset [Sun et al., 2021a], like CRIM13, consists of two mice interacting in an enclosed arena, captured by a top-view camera at 30 Hz. The dataset consists of videos with tracked poses and frame-level behavior annotations. Four behavior classes are labeled: attack, investigation, mount, and other (i.e., none of the above). The full training dataset consists of 506,668 frames across 68 videos; the test dataset consists of 262,107 frames across 19 videos. We accessed the pose estimates, TREBA features [Sun et al., 2021b], and behavior annotations from https://doi.org/10.22002/D1.1991 under the CC-BY 4.0 license. Frames for pretraining are available at https://doi.org/10.6084/m9.figshare.31337953 under the CC-BY 4.0 license.
@@ A.7 Facemap
Head-fixed mice were free to run on an air-floating ball in darkness [Syeda et al., 2024]. A single infrared camera captured one of several side or front views of the mouse’s face and upper trunk during each session at 50 Hz. Fifteen keypoints were labeled across the face (mouse, nose, whiskers, eyes) and paw. Neural activity was recorded across visual and sensorimotor areas using two-photon calcium imaging at 3 Hz. Approximately 30,000 to 50,000 cells were recorded in a given session. In our encoding task, we predict the 128 neural principal components following Syeda et al. [2024]. We evaluate neural encoding models on five randomly selected sessions. The publicly available data did not contain additional videos, so we only fine-tuned neural encoding models with this dataset. We accessed the raw videos, pose estimates, and neural activity from https://doi.org/10.25378/janelia.23712957 under the CC-BY-NC 4.0 license.
Table 3: Number of training/validation/test frames utilized across tasks, with number of source videos in parentheses. Pretraining frames are unlabeled, and the number includes both anchor frames and temporal context frames. Pose estimation and action segmentation frames are labeled; these models are trained across multiple videos. Neural encoding frames have matched neural activity (output) for each time point of behavior (input); these models are trained on single videos, since the neural populations change from one session to the next.
Pretraining Pose estimation Action segmentation Neural encoding train/val test train/val test train/val test IBL [IBL et al., 2025a] 270,025 (77) 7,609 (128) 1,446 (19) 35,521 (18) 7,786 (5) 338,760 (5) 42,720 (5) Mirror-mouse [Warren et al., 2021] 94,272 (17) 789 (17) 253 (5) - - - - Mirror-fish [Biderman et al., 2024] 47,943 (28) 373 (28) 94 (10) - - - - CRIM13 [Burgos-Artizzu et al., 2012] 99,926 (36) 3,986 (37) 1,274 (19) - - - - CalMS21 [Sun et al., 2021a] 91,087 (70) - - 506,668 (68) 262,107 (19) - - Facemap [Syeda et al., 2024] - - - - - 1,790,200 (5) 447,550 (5)
@@ Appendix B Beast implementation
Beast utilizes a standard ViT-B/16 architecture [Dosovitskiy et al., 2020], and combines a masked autoencoding and temporal contrastive learning loss. This approach is also taken by ViC-MAE [Hernandez et al., 2024], and we introduce key adaptations and simplifications to make Beast suitable for applications in behavioral neuroscience, which we elaborate on more in the following sections.